# TempoPFN: Synthetic Pre-training of Linear RNNs for Zero-shot Time Series Forecasting

### v2

@@ Appendix A Details on Synthetic Data Generation
CauKer. To increase the diversity and structural complexity of our training data, we used the CauKer generator from (Xie et al., 2024). This method produces multivariate time series by sampling from a structural causal model (SCM) where each variable is a Gaussian process. A random Directed Acyclic Graph (DAG) defines the dependencies between nodes, with each node having a maximum number of parents. Root nodes in the DAG are sampled from a GP prior $y_{i}\sim\mathcal{GP}(m(t),\kappa(t,t^{\prime}))$, using complex composite kernels $\kappa$ (combined via + or *) and stochastic mean functions $m(t)$ (e.g., linear $at+b$, exponential $a\exp(bt)$, or functions with anomalous impulses). Child nodes then apply nonlinear activation functions (e.g., ReLU, sigmoid, sin) to affine combinations of their parents’ values, introducing intricate, non-Gaussian dependencies.
@@
KernelSynth. The KernelSynth generator, based on Chronos (Ansari et al., 2024), samples independent univariate time series from Gaussian process priors $y\sim\mathcal{GP}(0,\kappa(t,t^{\prime}))$. It constructs composite kernels $\kappa$ by randomly combining base kernels (using addition or multiplication) from a large bank. This bank includes periodic kernels (ExpSineSquared($p$) with periods $p$ normalized by series length), stationary kernels (RBF, RationalQuadratic), and noise kernels (WhiteKernel). This method efficiently produces a vast array of smooth and structured series, ideal for learning fundamental temporal representations.
@@
Gaussian Process. The Gaussian Process generator, inspired by Mamba4Cast (Bhethanabhotla & Swelam, 2024), extends the GP sampling approach with greater complexity and realism. It constructs a composite kernel by combining up to six base kernels from a weighted bank that includes Matern, linear, periodic, and polynomial kernels. The combination logic (addition or multiplication) is also chosen randomly. To generate realistic periodicities, the periods of any periodic kernels are sampled from distributions tailored to the time series’ specified frequency (e.g., daily, weekly). Crucially, with a certain probability, we inject periodic peak spikes that are aligned with the dominant periodicity of the sampled kernel. This process creates sharp, recurring events on top of the smooth GP trajectory, yielding a wide range of both stationary and non-stationary series with complex covariance structures that mix smooth and abrupt dynamics.
@@
ForecastPFN. The ForecastPFN generator, adapted from Dooley et al. (2023), creates time series with configurable trends, seasonality, and noise patterns. The trend component combines linear and exponential elements multiplicatively for improved stability:
@@
The final series values are given by $\tau(t)\cdot s(t)\cdot(1+n(t))$, where $n(t)$ is Weibull-distributed noise. We enhanced this foundation with a noise injection strategy inspired by Bhethanabhotla & Swelam (2024), incorporating univariate augmentations like time warping, magnitude scaling, damping, and spike injection. A built-in filtering mechanism with retry logic ensures generated series avoid unrealistic spreads or extreme values, guaranteeing robust training data.
@@
A.1 Training Details and Hyperparameters
@@
Two-Stage Training Protocol. The training consists 4 million iterations with peak learning rate $2\times 10^{-4}$ and cosine annealing (Loshchilov & Hutter, 2017) schedulers (warmup ratio 0.003, minimum LR ratio 0.01).
We use the AdamW optimizer (Loshchilov & Hutter, 2019) with weight decay of 0.01, an effective batch size of approximately 200, and quantile loss over the [0.1–0.9] range. Other than weight decay, no explicit regularization such as dropout or early stopping is applied.
Architecture Selection. We ablated deeper models (8–16 layers) and found no consistent architectural winner. We therefore selected the 35M model (10 layers, embedding dimension 512, 4 heads) for its strong performance and computational efficiency.
@@
Model Total Parameters 34.69M
@@
Short convolution kernel size 16
@@
Training Total training series $\approx 10{,}000{,}000$
@@
Total training iterations 4,000,000
@@
Peak learning rate $2\times 10^{-4}$
@@
Appendix B Additional Experimental Details and Results
@@
Table 6: Ablation study of synthetic data priors where the model is trained on only a single prior at a time. To ensure tractability and manage computational costs, each ablation experiment was trained for a limited duration of 500,000 iterations. This should be contrasted with the several million iterations used for our final, fully-trained model, as detailed in the main paper. The ’Base Model’ was trained on all priors including augmentations. Lower values indicate better performance. Bold highlighting shows the best-performing individual priors, while underlined values indicate second-tier performance. Novel priors are our contributions, while Adapted priors are modified from open-source implementations.
@@
Table 7: Ablation study of core architectural components. All models were trained for 2M iterations with d=512, L=10, convolution size 16, and the base configuration uses H=4 Householder matrices with weaving enabled and negative eigenvalues allowed. We systematically ablate key components to understand their individual contributions. Results are sorted by overall CRPS (best to worst). Bold indicates best performance, underline indicates second-best. Lower values for CRPS and MASE indicate better performance.
@@
Table 8: Ablation study of model scale and depth. All models were trained for 4M iterations with H=4, convolution size 32, weaving enabled, and negative eigenvalues allowed. The Base Model uses d=512 embedding dimensions and L=10 layers. We explore the trade-off between model width (embedding dimension) and depth (number of layers) while keeping the parameter count roughly constant. Results are sorted by overall CRPS (best to worst). Bold indicates best performance, underline indicates second-best. Lower values for CRPS and MASE indicate better performance.
@@
Table 9: Ablation study of learning rate schedulers. All models use the base architecture (d=512, L=10, H=4, convolution size 32, weaving enabled) and were trained for 2M iterations with peak learning rate 2e-4. The WarmupStableDecay scheduler implements a three-phase approach with warmup (0.3% of steps), stable plateau (90% of steps), and cosine decay (9.7% of steps). CosineWithRestarts performs 4 periodic learning rate resets. CosineWithWarmup uses standard cosine annealing with warmup. Cosine applies basic cosine decay without warmup. Results are sorted by overall CRPS (best to worst). Bold indicates best performance, underline indicates second-best. Lower values for CRPS and MASE indicate better performance.
@@
B.1 Qualitative Comparison on the Gift-Eval Benchmark
This section presents qualitative examples of forecasts from the Gift-Eval benchmark. Each example is illustrated through a horizontal arrangement of four plots. The first plot on the left displays the full historical context. The subsequent three plots provide a zoomed-in view of the forecast horizon, comparing the predictions from top-performing zero-shot models in the following order: TempoPFN, TiRex, and TabPFN-TS.
It is important to note the different context lengths used by each model during evaluation. The TiRex baseline was evaluated using the full available context for each series. In contrast, TabPFN-TS utilized a context of 4096 time steps, while our model, TempoPFN, operated with a maximum context length of 3072.
@@ B.1 Qualitative Comparison on the Gift-Eval Benchmark
B.2 Quantitative Comparison on the Gift-Eval Benchmark
@@
Table 11: Normalized MASE scores of different zero-shot models on the GiftEval benchmark. Scores are relative to a seasonal naive baseline, where values below 1.0 indicate better performance. The models achieving the best and second-best scores are highlighted. )

### v3

@@ C.1 Generator Specifications
CauKer. To increase the diversity and structural complexity of our training data, we used the CauKer generator from (cauker). This method produces multivariate time series by sampling from a structural causal model (SCM) where each variable is a Gaussian process. A random Directed Acyclic Graph (DAG) defines the dependencies between nodes, with each node having a maximum number of parents. Root nodes in the DAG are sampled from a GP prior $y_{i}\sim\mathcal{GP}(m(t),\kappa(t,t^{\prime}))$, using complex composite kernels $\kappa$ (combined via + or *) and stochastic mean functions $m(t)$ (e.g., linear $at+b$, exponential $a\exp(bt)$, or functions with anomalous impulses). Child nodes then apply nonlinear activation functions (e.g., ReLU, sigmoid, sin) to affine combinations of their parents’ values, introducing intricate, non-Gaussian dependencies.
@@
KernelSynth. The KernelSynth generator, based on Chronos (ansari2024chronos), samples independent univariate time series from Gaussian process priors $y\sim\mathcal{GP}(0,\kappa(t,t^{\prime}))$. It constructs composite kernels $\kappa$ by randomly combining base kernels (using addition or multiplication) from a large bank. This bank includes periodic kernels (ExpSineSquared($p$) with periods $p$ normalized by series length), stationary kernels (RBF, RationalQuadratic), and noise kernels (WhiteKernel). This method efficiently produces a vast array of smooth and structured series, ideal for learning fundamental temporal representations.
@@
Gaussian Process. The Gaussian Process generator, inspired by Mamba4Cast (mamba4cast), extends the GP sampling approach with greater complexity and realism. It constructs a composite kernel by combining up to six base kernels from a weighted bank that includes Matern, linear, periodic, and polynomial kernels. The combination logic (addition or multiplication) is also chosen randomly. To generate realistic periodicities, the periods of any periodic kernels are sampled from distributions tailored to the time series’ specified frequency (e.g., daily, weekly). Crucially, with a certain probability, we inject periodic peak spikes that are aligned with the dominant periodicity of the sampled kernel. This process creates sharp, recurring events on top of the smooth GP trajectory, yielding a wide range of both stationary and non-stationary series with complex covariance structures that mix smooth and abrupt dynamics.
@@
ForecastPFN. The ForecastPFN generator, adapted from dooley2023forecastpfn, creates time series with configurable trends, seasonality, and noise patterns. The trend component combines linear and exponential elements multiplicatively for improved stability:
@@
The final series values are given by $\tau(t)\cdot s(t)\cdot(1+n(t))$, where $n(t)$ is Weibull-distributed noise. We enhanced this foundation with a noise injection strategy inspired by mamba4cast, incorporating univariate augmentations like time warping, magnitude scaling, damping, and spike injection. A built-in filtering mechanism with retry logic ensures generated series avoid unrealistic spreads or extreme values, guaranteeing robust training data.
@@
C.2 Synthetic Data Generation Throughput
In this section, we present the computational efficiency and resource flexibility of our pipeline. Unlike kernel-based methods such as KernelSynth, which can be computationally intensive due to the cubic $O(T^{3})$ complexity of Gaussian Processes, our approach enables high-throughput generation as shown in Table 6.
The benchmarking was conducted on a high-performance system featuring dual AMD EPYC 9334 32-Core Processors (128 threads total) and an NVIDIA L40S GPU. Crucially, the majority of our synthetic generators run exclusively on the CPU. The GPU is leveraged primarily for the few neural network-based prior models (e.g. Cauker).
Table 6: Profiling results for synthetic data generation throughput. N = Novel prior, A = Adapted from open-source.
Generator Source Length Series / Sec
Cauker A 2048 $0.66$
GP A 2048 $7.04$
Kernel A 2048 $0.32$
ForecastPFN A 2048 $35.49$
Sawtooth N 2048 $242.95$
Sinewave N 2048 $144.93$
Anomaly N 2048 $174.51$
Step N 2048 $106.58$
Stochastic Rhythm N 2048 $33.46$
Spike N 2048 $201.13$
SDE (OU Process) N 2048 $13.17$
Offline augmentations N 2048 $18.30$
Appendix D Training Details and Hyperparameters
@@
Training Protocol. For pretraining, we employ the AdamW optimizer (loshchilov-iclr19a) with a weight decay of 0.01 and batch size of 40. No additional regularization techniques, such as dropout or early stopping, are applied. Pretraining is conducted for 3 million iterations using a cosine annealing learning rate schedule (loshchilov-iclr17a) with a peak learning rate of $2\times 10^{-4}$, a warmup ratio of 0.003, and a minimum learning-rate ratio of 0.01. The model is trained using the quantile regression loss, computed independently for each output token across the set of quantile levels $\mathcal{Q}=\{q_{1},q_{2},\ldots,q_{m}\}$. In our experiments, we set $\mathcal{Q}=\{0.1,0.2,\ldots,0.9\}$ similarly as in TiRex and TabPFN-TS. The resulting losses are then averaged over all $h$ output tokens in a training sample. Given the true value $y_{t}$ at time $t$ and its predicted quantile value $\hat{y}^{(q)}_{t}$ for quantile level $q\in\mathcal{Q}$, the loss is defined as:
$L=\frac{1}{|\mathcal{Q}|\,h}\sum_{t=1}^{h}\sum_{q\in\mathcal{Q}}\begin{cases}q\,(y_{t}-\hat{y}^{(q)}_{t}),&\text{if }\hat{y}^{(q)}_{t}\leq y_{t},\\[6.0pt] (1-q)\,(\hat{y}^{(q)}_{t}-y_{t}),&\text{otherwise}.\end{cases}$
Architecture Selection. We ablated deeper models (8–16 layers) and found no consistent architectural winner. We selected the 10 layer model with embedding dimension of 512.
@@
Model Total Parameters 40M
@@
Short convolution kernel size 32
@@
Training Total training series 10M
@@
Total training iterations 3M
@@
Peak learning rate $7.5\times 10^{-3}$
@@
D.1 Additional Experimental Details and Results
@@
Table 8: Ablation study of single synthetic priors (trained for 500k iterations). ’Base Model’ uses all priors and augmentations. Lower values are better. Bold: best, underline: second-best. Novel priors are our contributions; Adapted are modified open-source versions.
@@
Table 9: Core architectural ablations (trained for 2M iterations with a per-GPU batch size of 40). Base config: $d=512$, $L=10$, conv size 16, $H=4$, weaving enabled, negative eigenvalues allowed. Sorted by overall CRPS. Bold: best, underline: second-best.
@@
Table 10: Ablation of model scale and depth (trained for 3M iterations with a batch size of 40). Base Model: $d=512,L=10$, $H=4$, conv size 32, weaving/neg eigenvalues on. Compares width vs. depth at constant parameter count. Sorted by overall CRPS. Bold: best, underline: second-best.
@@
Table 11: LR scheduler ablation (trained for 2M iterations with a batch size of 40). Base architecture: $d=512,L=10$, $H=4$, conv size 32, weaving enabled. WarmupStableDecay: warmup (0.3%), plateau (90%), cosine decay (9.7%). CosineWithRestarts: 4 resets. Sorted by overall CRPS. Bold: best, underline: second-best.
@@
Appendix E Comprehensive Quantitative Analysis
E.1 Computational Complexity and Efficiency
Table 12 summarizes the computational characteristics of TempoPFN relative to leading time-series foundation models, given sequence length $T$, horizon $H$, embedding dimension $d$, and layers $L$.
Training Complexity. Transformer-based models (Chronos, TimesFM, MOIRAI, TabPFN-TS) require $\mathcal{O}(T^{2}d)$ compute and memory due to self-attention, which becomes prohibitive for long context windows. TiRex reduces quadratic memory growth but remains sequential along $T$. In contrast, TempoPFN employs an associative GatedDeltaProduct recurrence, allowing parallel prefix-scan evaluation. This yields linear total work $\mathcal{O}(TLd^{2})$ and logarithmic parallel depth $\mathcal{O}(L\log T)$, enabling full sequence-length parallelism.
Inference Latency. Autoregressive models such as TiRex and Chronos must unroll $H$ steps to predict a horizon $H$, yielding $\mathcal{O}(H)$ latency. Transformer encoder models also scale their inference cost with $T^{2}$ even when used non-autoregressively. TempoPFN performs direct forecasting: concatenated query tokens allow the entire horizon to be produced in a single forward pass, giving constant $\mathcal{O}(1)$ latency with respect to $H$.
Table 12: Comparison of time-series foundation models. Complexities are reported with respect to sequence length $T$, horizon $H$, embedding dimension $d$, and layers $L$. We distinguish total work from parallel depth. TempoPFN benefits from scan-accelerated linear recurrences enabling sublinear depth.
Model Params (M) Training Time Inference Time Memory Parallelization
TiRex (auer2025tirex) 35
Work: $\mathcal{O}(TLd^{2})$
Depth: $\mathcal{O}(L\log T)$
AR: $\mathcal{O}(HLd^{2})$
Depth: $\mathcal{O}(L\log T)$
$\mathcal{O}(d)$ (streaming) Moderate (scan)
TabPFN-TS (hoo2024the) 11 $\mathcal{O}(T^{2}d)$ $\mathcal{O}(T^{2}d)$ $\mathcal{O}(Td)$ Moderate (attention)
TimesFM-2.0 (das2024decoder) 500 $\mathcal{O}(T^{2}d)$
Direct: $\mathcal{O}(T^{2}d)$
$\mathcal{O}(Td)$ High (transformer)
Chronos (ansari2024chronos) 9–205 $\mathcal{O}(T^{2}d)$
AR: $\mathcal{O}(H)$
$\mathcal{O}(Td)$ High (transformer)
MOIRAI-MoE (liu2024moirai) 14–935 $\mathcal{O}(T^{2}d)$ $\mathcal{O}(T^{2}d)$ $\mathcal{O}(Td)$ High (MoE + transformer)
TempoPFN (ours) 40
Work: $\mathcal{O}(TLd^{2})$
Depth: $\mathcal{O}(L\log T)$
Work: $\mathcal{O}(TLd^{2})$
Depth: $\mathcal{O}(L\log T)$
Horizon: $\mathcal{O}(1)$
$\mathcal{O}(d)$ (streaming)
or $\mathcal{O}(Td)$ (cached)
Full (sequence-parallel)
Figure 19: Comparison of TempoPFN normalized CRPS and MASE (3M iterations with a batch size of 40) against other models on Chronos-ZS benchmark.
Memory Usage. Transformer-based models require caching Key-Value pairs with $\mathcal{O}(Td)$ memory and, in some implementations, up to $\mathcal{O}(T^{2})$ activations. TiRex maintains a hidden state of size $\mathcal{O}(Td)$ during training. TempoPFN, being a Linear RNN, compresses the entire past into a single hidden state and supports streaming inference with constant $\mathcal{O}(d)$ memory, while still allowing optional $\mathcal{O}(Td)$ state storage if needed for analysis or hybrid decoding. Parallelization. Transformer-based models benefit from substantial batch-level parallelism but cannot eliminate the quadratic attention bottleneck. TiRex provides limited scan-style parallelism. TempoPFN achieves full sequence-level parallelization: the entire recurrence is computed via parallel scans, providing both high throughput and sublinear parallel depth. Overall, TempoPFN combines linear training cost, logarithmic parallel depth, constant-latency forecasting, and streaming memory usage, therefore, providing a zero-shot foundation model tailored for long-context forecasting settings.
E.2 Additional Results on Chronos-ZS and fev-bench
In this section, we provide additional empirical results on the Chronos-ZS and fev-bench benchmarks. On Chronos-ZS, Figure 19 shows the aggregated performance in terms of normalized probabilistic (CRPS) and point (MASE) forecasting, complementing Figure 7 in the main paper. On fev-bench, in addition to MASE (Table 2 in the main paper), we also present the leaderboard results for the Scaled Quantile Loss (SQL) in Table 13. SQL captures calibration quality by evaluating the quality of the entire predictive distribution at each time step. We can see that TempoPFN achieves Rank 6 again. In both metrics, our model outperforms the other leading synthetic-only baseline, TabPFN-TS (Rank 8 in both). Finally, to also visualize relative strengths in probabilistic forecasting, we show in Figure 20 the head-to-head Win Rates and Skill Scores based on SQL.
Table 13: fev-bench leaderboard based on Scaled Quantile Loss (SQL). Models are ranked by Win Rate and Skill Score.
Rank Model Avg. Win Rate (%) Skill Score (%) Median Runtime (s) Leakage (%) Failed Tasks (%) Organization Zero-shot
1 Chronos-2 91.3 47.3 3.57 0 0 AWS
2 TiRex 82.4 42.6 1.4 1 0 NX-AI
3 TimesFM 2.5 77.3 42.2 10.89 10 0 Google
4 Toto 1.0 69.9 40.7 77.51 8 0 Datadog
5 Moirai 2.0 63.6 39.3 1.9 28 0 Salesforce
6 TempoPFN 63.4 37.8 8.57 0 0 Uni Freiburg
7 Chronos-Bolt 63.2 38.9 1.0 0 0 AWS
8 TabPFN-TS 62.0 39.6 300.57 0 2 Prior Labs
9 Sundial-Base 44.4 33.4 33.99 1 0 Tsinghua University
10 Stat. Ensemble 43.8 20.2 624.45 0 11
11 AutoARIMA 39.0 20.6 120.16 0 10
12 AutoETS 32.6 -26.8 16.24 0 3
13 AutoTheta 25.9 5.5 9.27 0 0
14 Seasonal Naive 19.0 0.0 2.32 0 0
15 Naive 13.2 -45.4 2.24 0 0
16 Drift 9.0 -45.8 2.19 0 0
(a) Pairwise Win Rate (SQL) with 95% CIs
(b) Pairwise Skill Score (SQL) with 95% CIs
Figure 20: Head-to-head comparisons on FEV-Bench using Scaled Quantile Loss (SQL). (a) Win Rate: Percentage of tasks where the row model achieves lower error than the column model (ties count as half-wins); values >>50% indicate the row model is more accurate on average. (b) Skill Score: Average relative error reduction of the row model with respect to the column model; positive values indicate error reduction. Brackets indicate 95% confidence intervals estimated via 1000 bootstrap samples.
E.3 Feature-Space Alignment of Real and Synthetic Data Manifold
To empirically validate that our synthetic pre-training corpus effectively spans the manifold of real-world time series dynamics, we conducted a feature-space analysis comparing our synthetic data against the real-world benchmarks used for evaluation (GIFT-Eval, FEV-Bench, and Chronos).
Methodology. We randomly sampled up to 10,000 time series from each of our synthetic generators and the real-world datasets. For each series, we extracted a comprehensive vector of statistical time-series characteristics (including autocorrelation, approximate entropy, trend strength, spikiness, and seasonality metrics) using the tsfresh library (christ2018time). To visualize the relationship between these distributions, we standardized the feature vectors and projected them into a latent space using Uniform Manifold Approximation and Projection (UMAP) (mcinnes2018umap).
Analysis. Figure 21 presents the resulting embeddings in both 2D and 3D projections. The fact that real-world clusters are visible directly on top of the synthetic data confirms significant distributional overlap. The synthetic generators do not collapse into a single mode, but instead cover a vast region of the feature space, effectively “underpainting” the real-world benchmarks. This visual evidence supports our hypothesis that a diverse mixture of structurally distinct generators and data augmentations collectively covers the complex distribution of real-world temporal dynamics, enabling robust zero-shot transfer.
(a) 2D Projection ($N_{neigh}=5$, $d_{min}=0.1$, Silhouette Score: -0.054)
(b) 3D Projection ($N_{neigh}=5$, $d_{min}=0.4$, Silhouette Score: -0.001)
Figure 21: Feature-Space Distribution of Real vs. Synthetic Data. UMAP projections of time-series features extracted via tsfresh. Real-world benchmarks (GIFT-Eval, FEV-Bench, Chronos) are shown in cold colors (Blue/Teal) in the foreground, while our synthetic generators are shown in warm colors (Red/Orange/Yellow) in the background.
E.4 Qualitative Comparison on the Gift-Eval Benchmark
This section presents qualitative Gift-Eval forecasts in Figure 22, showing the full history (left) alongside zoomed-in predictions for TempoPFN, TiRex, and TabPFN-TS. Note that evaluation context lengths vary: TiRex uses the full history, TabPFN-TS uses 4096 steps, and TempoPFN uses 3072.
@@ E.4 Qualitative Comparison on the Gift-Eval Benchmark
E.5 Quantitative Comparison on the Gift-Eval Benchmark
@@
Table 15: Normalized MASE scores of different zero-shot models on the GiftEval benchmark. Scores are relative to a seasonal naive baseline, where values below 1.0 indicate better performance. The models achieving the best and second-best scores are highlighted.